mathslover (mathslover):

Solve the given differential equation. \[\cfrac{dy}{dx} = \cfrac{(x+y)^2}{(x+2)(y-2)}\]

mathslover (mathslover):

@ganeshie8 @Australopithecus @robtobey

OpenStudy (australopithecus):

What methods have you learned? I don't think this is separable

mathslover (mathslover):

I tried putting y = vx and x + y = v but none of them seem to work.

OpenStudy (australopithecus):

First bring everything to one side and make it equal to zero

mathslover (mathslover):

\(\cfrac{(x+2)(y-2) dy}{(x+y)^2 dx} - 1 = 0 \) What next?

OpenStudy (australopithecus):

That is in incorrect form

OpenStudy (australopithecus):

You want the form: \[M(x,y)dx + N(x,y)dy = 0\]

OpenStudy (australopithecus):

You can use substitution or you can check to see if it is exact and use that method

OpenStudy (australopithecus):

Substitution is as follows: You have two choices for substitution: u = x/y then, dx = udy + ydu Or u = y/x then, dy = xdu + udx

OpenStudy (australopithecus):

For setting it in the right form do you understand?

mathslover (mathslover):

Yeah, I've studied the method of substitution. In fact, I did try substituting y = vx but that didn't seem to work.

OpenStudy (australopithecus):

You wrote it in the wrong form so that could be your problem

mathslover (mathslover):

I understand that point. But, am not sure where to put that \((x+y)^2\) ... \((y-2)dy = \cfrac{dx}{x+2} \times (x+y)^2\)

OpenStudy (australopithecus):

So, for example: \[\frac{dx}{dy} = \frac{(y + x)}{(x + 2y)^2}\] in standard form would be: \[dx(x+2y)^2 - (y+x)dy = 0\] M(x,y) = (x + 2y)^2 N(x,y) = -(y+x)

OpenStudy (australopithecus):

Sorry this would be standard form \[(x+2y)^2dx - (y+x)dy = 0\]

mathslover (mathslover):

Okay, so, I have : \((y-2)(x+2) dy - (x+y)^2 dx = 0 \) This is the standard form, right?

OpenStudy (australopithecus):

Yup :D

OpenStudy (australopithecus):

Ok now you have to test to see if you can even use the separation method

OpenStudy (australopithecus):

Or substitution sorry

OpenStudy (australopithecus):

This is clearly not seperable

mathslover (mathslover):

Now, I will simply use substitution : y = vx \(dy = vdx + xdv \) The equation becomes: \((vx - 2)(x+2)(vdx + xdv) -(x+vx)^2 dx = 0 \) \((vx-2)(x+2)(vdx + xdv) = x^2 (v+1)^2 dx \) Should I proceed further or am I doing something wrong?

OpenStudy (australopithecus):

Expand (y-2)(x+2)

OpenStudy (australopithecus):

Also expand (x+y)^2

OpenStudy (australopithecus):

MY browser is lagging like crazy one second

OpenStudy (australopithecus):

there is a test you have to do to make sure you can even use substitution method

mathslover (mathslover):

Okay, it will give me : \[(y-2)(x+2) dy - (x+y)^2 dx = 0 \\ xydy + 2(y-x) dy - 4dy - x^2 dx -y^2 dx -2xy dx = 0 \\ xydy - 2y dy - 2xdy - x^2 dx - y^2 dx - 2xydx = 0 \]

mathslover (mathslover):

Test? Never heard of it. Can you please explain?

OpenStudy (australopithecus):

Now factor out dy and dx

mathslover (mathslover):

\[(xy - 2y - 2x)dy - (x^2 + y^2 + 2xy)dx = 0 \\ (xy - 2y - 2x)dy = (x^2 + y^2 + 2xy)dx \] Seems like I'm reverting everything back to the original equation :/ :(

OpenStudy (australopithecus):

You can only do substitution method if \[M(\lambda x, \lambda y) = \lambda^pM(x,y)\ for\ all\ \lambda,x,y\] and \[N(\lambda x, \lambda y) = \lambda^qN(x,y)\ for\ all\ \lambda,x,y\] to use separation p must equal q

OpenStudy (australopithecus):

First to check if M(x,y) or N(x,y) is homogenous here are some examples:

OpenStudy (australopithecus):

The degree of homogeneity must be equal for both terms N(x,y) and M(x,y) to use the substitution method

OpenStudy (australopithecus):

hence p = q

OpenStudy (australopithecus):

its a pretty simple check you can do in your mind almost

mathslover (mathslover):

I seem to get your point! Will give it a try next morning. Gotta sleep now! :) Thanks a lot for your help! I'll let you know if I get stuck anywhere.

OpenStudy (australopithecus):

Wait lets at least check to see

OpenStudy (australopithecus):

because you might have to use another method

mathslover (mathslover):

Yeah sure!

OpenStudy (australopithecus):

Is M(x,y) homogenous?

OpenStudy (australopithecus):

and to what degree

OpenStudy (australopithecus):

Sorry I know you are probably tired just I want to save you time

OpenStudy (australopithecus):

If you are lost I dont mind giving you another example

mathslover (mathslover):

It doesn't seem to be homogeneous :( because of that constant there

OpenStudy (australopithecus):

So you have to use a different method

mathslover (mathslover):

Oh! :( Any ideas for which method to use?

OpenStudy (australopithecus):

Yup

OpenStudy (australopithecus):

You can check to see if it is exact

OpenStudy (australopithecus):

Do you want to sleep we can work on this later if you want?

OpenStudy (australopithecus):

its actually a pretty long method

OpenStudy (australopithecus):

kind of

mathslover (mathslover):

Oh! Well, I will love to give it a try now. I can control my sleep for studies!

OpenStudy (australopithecus):

Alright, first step you need to test to see if the DE is exact or not

OpenStudy (australopithecus):

A DE is said to be exact when, \[M_y(x,y) = N_x(x,y)\] You know how to take partial derivatives right?

mathslover (mathslover):

Yeah!

OpenStudy (australopithecus):

Also I would put it in another form but I am not sure how to use write daran in latex

OpenStudy (australopithecus):

So first step take both partial derivatives and check to see if they are equal

OpenStudy (australopithecus):

if they are not this question is even longer ha

mathslover (mathslover):

\(\delta\) You mean this?

OpenStudy (australopithecus):

that is lower case delta

mathslover (mathslover):

\(\Delta\) .. ?

OpenStudy (australopithecus):